A new car that runs on electricity and gasoline is advertised as getting 65 mpg with a standard deviation of 6.9. If you test drive a car, what is the probability that it will get between 50 and 65 mpg?
Answer by stanbon(75887)

A new car that runs on electricity and gasoline is advertised as getting 65 mpg with a standard deviation of 6.9. If you test drive a car, what is the probability that it will get between 50 and 65 mpg?
-------------
z(50) = (50-65)/6.9 = -15/6.9 = -2.1739
z(65) = 0
----
P(50 < x < 65) = P(-2.1739 < z < 0) = normalcdf(-2.1739,0) = 0.4851
--------------
Cheers,
Stan H.
----------------